# RAIFLE: Reconstruction Attacks on Interaction-based Federated Learning with Adversarial Data Manipulation

### v2

@@ I Introduction
In light of these findings, we discuss the pros and cons of various potential countermeasures to alleviate the risks of our privacy attacks, including local differential privacy, secure aggregation, data validation, personalization, etc. We hope our research can inform future endeavors to develop more secure and private federated RS/OLTR systems specifically and IFL in general. To ensure the reproducibility of our results, we have made our code publicly available at https://github.com/dzungvpham/raifle.
To summarize, our main contributions are:
@@
Section II provides the background and related work. Section III and IV present RAIFLE and various implementation details. Section V and VI present our evaluation setup and results. Section VII analyzes countermeasures against RAIFLE. Section VIII discusses RAIFLE’s limitations and next steps.
@@ III-C3 Indirect Manipulation
In certain scenarios, the FL server cannot arbitrarily alter the training feature values however it wants. While lacking direct control, the server can still indirectly influence the extracted features by modifying the representation of the items. We focus on the case where the features of the items are extracted locally by the users using a fixed procedure agreed upon by all participants. For example, in an image recommendation system, the server can only present images to users, and the image features can only be extracted using some pre-trained computer vision model for object classification. However, the server can still execute the noise injection method by adding small perturbations to the images such that the extracted features closely resemble the target noise. The success of this method highly depends on how pliable the feature extractor is. Our experiments with image-based ranking show that existing computer vision models are highly susceptible to our adversarial manipulation, which leads to improved reconstruction performance (Section VI-C). We further enhance the reconstruction success by integrating the fingerprinting technique (Section IV-C).
To summarize, we present the high-level pseudocode for RAIFLE with ADM in Algorithm 1. Details about specific ADM implementations can be found in Section III-C2 and V.
@@ IV Implementation Details
Pointwise: If $\mathcal{L}_{FL}$ is not differentiable w.r.t. to $\mathcal{I^{\prime}}$ (e.g. 0-1 classification loss), change it to a differentiable one such as $L_{2}$ loss. In fact, if $\mathcal{L}_{FL}$ is $L_{2}$ loss, then we can show that $\nabla_{\mathcal{I^{\prime}}}^{2}g=\mathbf{0}$ (see Appendix -A), which implies that the attack is convex (Theorem 1).
@@ VI Evaluation Results
In this section, we present our evaluation results for the scenarios described in Section V. (More detailed results can be found in Appendix -C.) We primarily report the area under the curve of the receiver operating characteristic (AUC) [15] since we do not want to rely on a fixed threshold for the classification and we equally care about positive and negative interactions. Note that randomly guessing will result in an AUC of $\approx 0.5$. We summarize some highlights of our findings as follows:
@@ VII Countermeasures
In this section, we look at several possible countermeasures for RAIFLE, including local differential privacy (LDP), secure aggregation (SA), manipulation detection, and other techniques.
@@ VII-A Local Differential Privacy
To test the effect of LDP on RAIFLE, we experiment with the Gaussian mechanism by applying Gaussian noise to each user’s local update using privacy budget $\varepsilon\in\{1,20,100,500\}$ ($\delta=10^{-8}$). Table IX shows the results of applying LDP on two example configurations from each evaluation scenario in Section V (FNCF with ML-100K and Steam-200K for federated RS; FPDGD with linear ranker, informational click model, and 16 queries on MQ 2007 and MSLR-WEB10K; neural ranker with 8 hidden units and number of items equal to number of features using ResNet18 and DenseNet121). $\varepsilon$ is chosen to represent a wide range of privacy protections, with $\varepsilon=1$ providing particularly strong guarantees under LDP (at the expense of utility) and $\varepsilon=500$ representing a typical LDP budget adopted in LDP research and applications [10, 27]. We set the sensitivity $\Delta_{2}=0.1,0.5,0.05$ for federated RS, FPDGD, and FOLTR with images, respectively. We use the Analytical Gaussian formula [7] for $\varepsilon>1$ and the standard one for $\varepsilon=1$ [20]. For more LDP results, see Appendix -C.
@@
TABLE X: FPDGD’s average test NDCG@10 vs LDP on MSLR-10K.
@@
Despite the capability of LDP, it can significantly reduce the utility of the learned FL model [70]. To illustrate the privacy-utility tradeoff, we simulate the FL training for the FPDGD scenario on the MSLR-WEB10K dataset with a linear model and a neural net with 16 hidden units. The FL model is trained on a single pass of the train portion in Fold 1 of the dataset, where each user is assigned one unique query. FL aggregation occurs for every 100 users. We measure the normalized discounted cumulative gain for the top 10 items (NDCG@10) on the corresponding test portion. The simulation is repeated 10 times. Our results show that even for $\varepsilon\!=\!100$, the FL model’s performance suffers a noticeable decrease (Table X). Thus, in reality, FL practitioners often need to use large LDP $\varepsilon$ in the tens or hundreds range to get meaningful model performance [10, 27]. Determining the right balance between privacy protection and application utility is essential to adequately defend against attacks like RAIFLE. For more results on the privacy-utility tradeoff of FNCF and FPDGD, we refer readers to Table 4 in the IMIA paper [74] for FNCF and Table 2 in the FPDGD paper [67].
@@ VII-B Secure Aggregation with DP
VII-C Private Information Retrieval
VII-D Detecting Data Manipulation
VII-D1 Cryptography
To prevent the server from executing ADM, users can validate the integrity and authenticity of the items sent by the server using cryptography techniques. For example, users can employ cryptographic hashes or digital signatures to ensure that their training items are the same as all of their peers (similar to the parameters validation defense for model manipulation in [47]). However, different users can have different items, which renders the process of cross-checking items more difficult, not to mention the need to do so in a privacy-preserving manner to avoid leaking information to other users. Furthermore, detecting inconsistency among users would fail if the server manipulates the data in the same manner for all users.
VII-D2 Heuristic Checks
@@
Another method is to directly check the training data for any sign of ADM. Detecting our fingerprinting method is straightforward: simply checking for the presence of zeroed-out (or constant) features would suffice. Detecting noise injection, however, is more difficult: while we can try to quantify the degree of randomness in the data (e.g., via Shannon entropy), a malicious server can mask the noise, such as by mixing noise and real data via a convex combination. As another example, our image-based attack (Section IV-C) can include a regularizer to further reduce the visual artifacts (Appendix -B). Designing a comprehensive set of data quality checks is likely not practically feasible due to the vast possibilities of manipulation techniques and data modalities. The difficulty of detecting active manipulation is also reflected in model manipulation attacks [12].
VII-E Minimizing Shared Information
@@ VIII-A Impact on FL Utility
While RAIFLE requires some modifications to the local learning algorithm (Section III-B), this modified version is only used by the server for performing the gradient matching optimization and is never used by real users, thus never affecting the FL model. The use of manipulated data for ADM, however, will certainly reduce the performance of the global FL model if not handled properly. Depending on the FL setup and the amount of control available, the server can take specific measures to limit the negative impacts of ADM. If SA is not applied, the server can simply disregard all FL updates from targeted users since their identities are known. Should discarding bad FL updates be infeasible, the server can choose a small number of users to apply ADM in one or a few FL rounds only instead of attacking every user in every round, thereby minimizing the low-quality contributions from targeted users. If SA is applied, ADM will have to be performed for every user (Section VII-B), thus requiring the server to discard or significantly reduce the weight of the affected FL rounds’ results, which is achievable if the server can control each FL round and the participants.
VIII-B Relationship to Model Manipulation and Other Attacks
@@
Aside from model manipulation, RAIFLE also shares some minor similarities with other ML attacks while still exhibiting noteworthy differences. RAIFLE has a similar goal to that of gradient inversion [76, 46] but focuses more on inverting the “labels” produced by the users. It involves modifying input data like the adversarial perturbation or backdoor attacks [3, 6, 38], although its aim is not to mislead models but to extract more information from users. RAIFLE thus can be considered a distinct blend of existing classes of ML attacks designed for the purpose of attacking IFL or similar scenarios. We believe our attack set a novel research direction in ML system vulnerabilities and exploits, particularly concerning interactive ML systems where users do not have full control over their experiences.
VIII-C Using RAIFLE as Guardrail
VIII-D Limitations and Future Directions
In this work, we mainly use numeric representations of the items that users interact with. Our method is not yet extensible to non-differentiable item representations such as natural text. User interactions are also currently restricted to values to which the reconstruction can be made differentiable. As such, we intend to develop ADM techniques for more types of item modalities, including multimodal representations (e.g., image and text), and to investigate more IFL scenarios in which the interactions cannot be easily “smoothened”. Our proposed ADM techniques involve manipulating data only and were developed with knowledge of the underlying FL algorithm. It is potentially possible to combine data manipulation with model manipulation [21, 47, 12] to achieve a better attack. Automatic discovery of ADM, particularly for determining the best noise injection method, is also of special interest. Lastly, we intend to look into more covert ADM methods by introducing a manipulation budget while retaining strong reconstruction performance. Our preliminary results in this direction via restricting the number of manipulated features show a rather noticeable reduction in the attack’s capability, which we hope to remediate. We leave these for future research.
@@ Acknowledgment
We thank Professor Negin Rahimi for her helpful comments on learning-to-rank and information retrieval.

### v3

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 24-28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.240363 www.ndss-symposium.org
@@ I Introduction
In light of these findings, we discuss the pros and cons of various potential countermeasures to alleviate the risks of our privacy attacks, including local differential privacy, secure aggregation, data validation, personalization, etc. We hope our research can inform future endeavors to develop more secure and private federated RS/OLTR systems specifically and IFL in general. To summarize, our main contributions are:
@@ III-C3 Indirect Manipulation
In certain scenarios, the FL server cannot arbitrarily alter the training feature values however it wants. While lacking direct control, the server can still indirectly influence the extracted features by modifying the representation of the items. We focus on the case where the features of the items are extracted locally by the users using a fixed procedure agreed upon by all participants. For example, in an image recommendation system, the server can only present images to users, and the image features can only be extracted using some pre-trained computer vision model for object classification. However, the server can still execute the noise injection method by adding small perturbations to the images such that the extracted features closely resemble the target noise [65]. The success of this method highly depends on how pliable the feature extractor is. Our experiments with image-based ranking show that existing computer vision models are highly susceptible to our adversarial manipulation, which leads to improved reconstruction performance (Section VI-C). We further enhance the reconstruction success by integrating the fingerprinting technique (Section IV-C).
To summarize, we present the high-level pseudocode for RAIFLE with ADM in Algorithm 1. Details about specific ADM implementations can be found in Section IV-C and V.
@@ IV Implementation Details
Pointwise: If $\mathcal{L}_{FL}$ is not differentiable w.r.t. to $\mathcal{I^{\prime}}$ (e.g. 0-1 classification loss), change it to a differentiable one such as $L_{2}$ loss. In fact, if $\mathcal{L}_{FL}$ is $L_{2}$ loss, then we can show that $\nabla_{\mathcal{I^{\prime}}}^{2}g=\mathbf{0}$ (see Appendix A-A), which implies that the attack is convex (Theorem 1).
@@ VI Evaluation Results
In this section, we present our evaluation results for the scenarios described in Section V. We primarily report the area under the curve of the receiver operating characteristic (AUC) [16] since we do not want to rely on a fixed threshold for the classification and we equally care about positive and negative interactions. Note that randomly guessing will result in an AUC of $\approx 0.5$. We summarize some highlights of our findings as follows:
@@ VII Countermeasures
In this section, we look at several possible countermeasures for RAIFLE, including local differential privacy (LDP) and secure aggregation (SA). See Appendix A-C for more discussion.
@@ VII-A Local Differential Privacy
To test the effect of LDP on RAIFLE, we experiment with the Gaussian mechanism by applying Gaussian noise to each user’s local update using privacy budget $\varepsilon\in\{1,20,100,500\}$ ($\delta=10^{-8}$). Table IX shows the results of applying LDP on two example configurations from each evaluation scenario in Section V (FNCF with ML-100K and Steam-200K for federated RS; FPDGD with linear ranker, informational click model, and 16 queries on MQ 2007 and MSLR-WEB10K; neural ranker with 8 hidden units and number of items equal to number of features using ResNet18 and DenseNet121). $\varepsilon$ is chosen to represent a wide range of privacy protections, with $\varepsilon=1$ providing particularly strong guarantees under LDP (at the expense of utility) and $\varepsilon=500$ representing a typical LDP budget adopted in LDP research and applications [11, 28]. We set the sensitivity $\Delta_{2}=0.1,0.5,0.05$ for federated RS, FPDGD, and FOLTR with images, respectively. We use the Analytical Gaussian formula [8] for $\varepsilon>1$ and the standard one for $\varepsilon=1$ [21].
@@
TABLE X: FPDGD’s avg. test NDCG@10 vs LDP $\varepsilon$ (MSLR-10K).
@@
Despite the capability of LDP, it can significantly reduce the utility of the learned FL model [78]. To illustrate the privacy-utility tradeoff, we simulate the FL training for the FPDGD scenario on the MSLR-WEB10K dataset with a linear model and a neural net with 16 hidden units. The FL model is trained on a single pass of the train portion in Fold 1 of the dataset, where each user is assigned one unique query. FL aggregation occurs for every 100 users. We measure the normalized discounted cumulative gain for the top 10 items (NDCG@10) on the corresponding test portion. The simulation is repeated 10 times. Our results show that even for $\varepsilon\!=\!100$, the FL model’s performance suffers a noticeable decrease (Table X). Thus, FL practitioners often need to use large LDP $\varepsilon>10$ to achieve practical model performance [11, 28]. Determining the right balance between privacy protection and application utility is essential to adequately defend against attacks like RAIFLE. For more results on the privacy-utility tradeoff of FNCF and FPDGD, we refer readers to Table 4 in the IMIA paper [82] and Table 2 in the FPDGD paper [74].
@@ VII-B Secure Aggregation with DP
VII-C Detecting Data Manipulation
VII-C1 Cryptography
To prevent the server from executing ADM, users can validate the integrity and authenticity of the items sent by the server using cryptography techniques. For example, users can compare cryptographic hashes (i.e., checksum) of the items to ensure that their training features are the same as all of their peers (similar to the parameters validation defense for model manipulation in [53]). However, different users can have different items, which renders the process of cross-checking items more difficult, not to mention the need to do so in a privacy-preserving manner to avoid leaking information to other users. Furthermore, detecting inconsistency among users would fail if the server manipulates the data in the same manner for all users.
VII-C2 Heuristic Checks
@@
Another method is to directly check the training data for any sign of ADM. Detecting our fingerprinting method is straightforward: simply checking for the presence of zeroed-out (or constant) features would suffice. Detecting noise injection, however, is more difficult: while we can try to quantify the degree of randomness in the data (e.g., via Shannon entropy), a malicious server can mask the noise, such as by mixing noise and real data via a convex combination. As another example, our image-based attack (Section IV-C) can include a regularizer to further reduce the visual artifacts (Appendix A-B). Designing a comprehensive set of data quality checks is likely not practically feasible due to the vast possibilities of manipulation techniques and data modalities. The difficulty of detecting active manipulation is also reflected in model manipulation attacks [13].
VII-D Minimizing Shared Information
@@ VIII-A Impact on FL Utility
While RAIFLE requires some modifications to the local learning algorithm (Section III-B), this modified version is only used by the server for performing the gradient matching optimization and is never used by real users, thus never affecting the FL model. The use of manipulated data for ADM, however, will certainly reduce the performance of the global FL model if not handled properly. Depending on the FL setup and the amount of control available, the server can take specific measures to limit the negative impacts of ADM. If SA is not applied, the server can simply disregard all FL updates from targeted users since their identities are known. Should discarding bad FL updates be infeasible, the server can choose a small number of users to apply ADM in one or a few FL rounds only instead of attacking every user in every round, thereby minimizing the low-quality contributions from targeted users. As a demonstration, from Figure 7, the test NDCG@10 for a linear FPDGD ranker trained on the MSLR-WEB10K dataset with 100 users is only reduced by $\approx$ 5-10% even when 20-40% of users have random item features. If SA is applied, ADM will have to be performed for every user (Section VII-B), thus requiring the server to discard or significantly reduce the weight of the affected FL rounds’ results, which is achievable if the server can control each FL round and the participants.
Figure 7: Utility (test NDCG@10) vs number of attacked users (out of 100) for a linear FPDGD ranker on the MSLR-WEB10K dataset.
VIII-B Constrained Server Capabilities
TABLE XII: Percentage of manipulated features vs Mean reconstruction AUC for FPDGD on MQ2007 and MSLR-WEB10K
Dataset Model Informational Navigational
0% 50% 75% 100% 0% 50% 75% 100%
MQ2007 (16 queries) Linear 0.65 0.77 0.79 0.82 0.71 0.86 0.87 0.88
Neural 4 0.57 0.72 0.88 0.95 0.60 0.79 0.95 0.98
Neural 8 0.56 0.71 0.91 0.99 0.58 0.78 0.98 1.00
Neural 16 0.53 0.67 0.90 1.00 0.55 0.74 0.98 1.00
MSLR-10K (48 queries) Linear 0.50 0.51 0.65 0.80 0.50 0.52 0.70 0.86
Neural 4 0.50 0.50 0.54 0.87 0.50 0.50 0.55 0.93
Neural 8 0.50 0.50 0.52 0.91 0.50 0.50 0.53 0.96
Neural 16 0.50 0.50 0.51 0.91 0.51 0.50 0.51 0.96
In certain settings, the central server might be limited in its capability to manipulate the training features or even be non-existent (as in the case of decentralized federated recommendation systems; see Appendix A-C3). Here, we consider a constrained scenario where the server can only control a subset of the features. We repeat the FPDGD experiment (Section V-B) but restrict to manipulating only 50% and 75% of all features. Our results from the hardest settings tested for both MQ2007 and MSLR-WEB10K (without DP) indicate a positive correlation between RAIFLE’s attack performance and the percentage of manipulated features (Table XII). However, we can observe that the impact varies depending on the dataset and the number of items, with MSLR-WEB10K incurring a much steeper reduction in mean attack AUC compared to MQ2007, essentially no better than random guessing when manipulating only 50% of features. We can also see that the effect of increasing the number of model parameters on attack AUC is mostly reversed when manipulating $<100$% of features, with more features leading to weaker reconstruction performance (except for MQ2007 at 75% manipulation). In the case where no features can be manipulated but the presentation of the items can be modified (e.g., ordering), it might be possible for the server to exploit users’ click bias w.r.t. the item positions or popularity [82]. However, it is unclear how to turn this angle into an effective or efficient attack since we cannot readily leverage such discrete manipulation for optimization.
VIII-C Computational and Storage Overhead
In our experiments, image-based manipulation is the only scenario where our attack needs significantly more computational resources. With a single Nvidia Tesla V100 16GB GPU, it takes $\approx$ 5 minutes to manipulate each batch of 256 images for DenseNet121, the largest model tested. Note that this step only needs to be done once since the manipulated images can be reused for different users. Performing the gradient inversion in this scenario is much faster: only $\approx$ 20 seconds for each user with 4096 images on DenseNet121. Considering the massive computing capabilities of adversaries like Google or Facebook, our attack is very cheap to perform, not to mention it can be executed separately from the FL protocol, thus avoiding any impact on the FL protocol’s speed. With regards to storage space, in the case of images, we would need to additionally store at least one extra manipulated image for every original image. However, in the case of direct non-image manipulation, we can simply store a single random seed for each user so that we can recreate the attack noise vector anytime.
VIII-D Relationship to Model Manipulation and Other Attacks
@@
Aside from model manipulation, RAIFLE also shares some minor similarities with other ML attacks while still exhibiting noteworthy differences. RAIFLE has a similar goal to that of gradient inversion [84, 52] but focuses more on inverting the “labels” produced by the users. It involves modifying input data like the adversarial perturbation or backdoor attacks [3, 7, 41], although its aim is not to mislead models but to extract more information from users. RAIFLE thus can be considered a distinct blend of existing classes of ML attacks designed to attack IFL or similar scenarios. We believe our attack set a novel research direction in ML system vulnerabilities and exploits, particularly interactive ML systems where users do not have full control over their experiences.
VIII-E Limitations and Future Work
RAIFLE is currently restricted to FL settings wherein a central server controls the features of user interaction items, with federated recommendation and learning-to-rank as the two most prominent examples. In this sense, RAIFLE is not as flexible as model manipulation attacks [22, 53, 13] which can be applied in more general (but still malicious) FL scenarios. However, we argue that federated RS and OLTR are among the most important applications of FL to study given how widely-used services like search engines, e-commerce product suggestions, and entertainment recommendations all collect a vast amount of sensitive user data. Furthermore, there have been several prominent real-world cases of large-scale recommendation data manipulation, with companies like Facebook, Amazon, and Google manipulating the new feeds or search recommendations for hundreds of thousands of users [50, 62, 5]. These examples highlight the relevance of our strong threat model and the genuine risk of centralized entities exploiting their positions for their own benefit.
With regards to RAIFLE’s design, this work mainly explores numeric feature representations only and is not yet extensible to non-differentiable item representations such as natural text. User interactions are also currently restricted to values to which the reconstruction can be made differentiable. As such, we intend to develop ADM techniques for more types of item modalities, including multimodal representations (e.g., image and text), and to investigate more IFL scenarios in which the interactions cannot be easily “smoothened”. Our proposed ADM techniques involve manipulating data only and were developed with knowledge of the underlying FL algorithm. It is potentially possible to combine data manipulation with model manipulation [22, 53, 13] to achieve a better attack. Automatic discovery of ADM, particularly for determining the best noise injection method, is also of special interest. Lastly, we intend to look into more covert ADM methods by introducing a manipulation budget while retaining strong reconstruction performance. Our preliminary results in this direction via restricting the number of manipulated features show a rather noticeable reduction in the attack’s capability, which we hope to remediate. We leave these for future research.
@@ Acknowledgment
We would like to thank Professor Negin Rahimi for her helpful comments on learning-to-rank and information retrieval. The project was supported in part by NSF grant 2131910.
Appendix A
A-A Closed-form Solution for RAIFLE
We present below the derivation for a closed-form solution for $\mathcal{I^{\prime}}$ under certain assumptions regarding the learning algorithm $g$.
We are given items $\mathcal{X}$, a model $f$ with global model parameters $\theta$, updated model parameters $\hat{\theta}$, and learning algorithm $g$. Suppose that $\mathcal{L}_{atk}$ is the $L_{2}$ loss and $\mathcal{L}_{FL}$ is the pointwise $L_{2}$. We further assume that:
$g(\mathcal{X},\theta,\mathcal{I}^{\prime})$ $=\nabla_{\theta}\mathcal{L}_{FL}(\mathcal{I^{\prime}},f(\mathcal{X},\theta))$ (8)
$=\nabla_{\theta}||\mathcal{I^{\prime}}-f(\mathcal{X},\theta)||^{2}_{2}$
$=-2\nabla_{\theta}f(\mathcal{X},\theta)\cdot(\mathcal{I^{\prime}}-f(\mathcal{X},\theta))$
Thus, we have:
$\nabla_{\mathcal{I^{\prime}}}g(\mathcal{X},\theta,\mathcal{I}^{\prime})=-2\nabla_{\theta}^{T}f(\mathcal{X},\theta)$ (9)
and consequently:
$\nabla^{2}_{\mathcal{I^{\prime}}}g(\mathcal{X},\theta,\mathcal{I}^{\prime})=\mathbf{0}$ (10)
Therefore, by theorem 1, there exists a global optimum for RAIFLE. Setting $\nabla_{\mathcal{I^{\prime}}}\mathcal{L}_{atk}(\hat{\theta},\theta^{\prime})$ to $\mathbf{0}$ (from eq. 6), we have:
$\nabla_{\mathcal{I^{\prime}}}\mathcal{L}_{atk}(\hat{\theta},\theta^{\prime})=\mathbf{0}$ (11)
$\iff$ $\nabla_{\mathcal{I^{\prime}}}g(\mathcal{X},\theta,\mathcal{I^{\prime}})\cdot(\hat{\theta}-g(\mathcal{X},\theta,\mathcal{I^{\prime}}))=\mathbf{0}$
$\iff$ $\nabla_{\mathcal{I^{\prime}}}g(\mathcal{X},\theta,\mathcal{I^{\prime}})\cdot\hat{\theta}=\nabla_{\mathcal{I^{\prime}}}g(\mathcal{X},\theta,\mathcal{I^{\prime}})\cdot g(\mathcal{X},\theta,\mathcal{I^{\prime}})$
Substituting the results from eq. 8 and 9 into eq. 11, we can now derive the closed-form solution for $\mathcal{I^{\prime}}$:
$\nabla_{\mathcal{I^{\prime}}}\mathcal{L}_{atk}(\hat{\theta},\theta^{\prime})=\mathbf{0}$ (12)
$\Leftrightarrow$ $\nabla_{\theta}^{T}f(\mathcal{X},\theta)\cdot\hat{\theta}=-2\nabla_{\theta}^{T}f(\mathcal{X},\theta)\nabla_{\theta}f(\mathcal{X},\theta)\cdot(\mathcal{I^{\prime}}-f(\mathcal{X},\theta))$
$\Leftrightarrow$ $\mathcal{I^{\prime}}=f(\mathcal{X},\theta)-\frac{1}{2}\left(\nabla_{\theta}^{T}f(\mathcal{X},\theta)\nabla_{\theta}f(\mathcal{X},\theta)\right)^{-1}\nabla_{\theta}^{T}f(\mathcal{X},\theta)\cdot\hat{\theta}$
Note that $\nabla_{\theta}^{T}f(\mathcal{X},\theta)\nabla_{\theta}f(\mathcal{X},\theta)$ is invertible iff $\operatorname{\mathbf{rank}}(\nabla_{\theta}f(\mathcal{X},\theta))=m$. Therefore, eq. 12 is a valid closed-form solution if and only if $\operatorname{\mathbf{rank}}(\nabla_{\theta}f(\mathcal{X},\theta))=m$.
A-B Indirect Manipulation’s Image Quality
To measure the similarity of the manipulated images to the originals in our federated OLTR experiment with image-based data (Section V-C), we calculated the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM) [76]. Overall, the average SSIM ranges from 0.70 to 0.80 and the average PSNR ranges from 25 to 30 across all tested computer vision models (Table XIII). These values indicate that the adversarial images are relatively similar to the originals but not without visible artifacts or distortions. While our research is not focused on creating the most subtle manipulations, it is straightforward to include a regularization loss based on SSIM or PSNR in the manipulation process to improve image similarity, likely at the expense of attack performance. We leave this for future research.
TABLE XIII: SSIM and PSNR of manipulated images and originals (rounded to 2nd decimal place)
Model SSIM Mean SSIM Std. PSNR Mean PSNR Std.
ResNet18 0.70 0.06 25.80 1.39
RegNet Y 800MF 0.80 0.06 29.36 1.27
DenseNet121 0.77 0.05 27.60 1.67
MNasNet 1.3 0.76 0.06 28.00 1.30
A-C Discussion of Additional Countermeasures
We briefly discuss several hypothetical defense mechanisms against RAIFLE, including PIR, self-attacking, and alternative FL protocols. While these approaches are not yet as practical as the ones discussed in Section VII, they are nonetheless theoretically interesting to consider.
A-C1 Private Information Retrieval
A-C2 Using RAIFLE as Guardrail
A-C3 Alternative FL Protocols
RAIFLE’s threat model assumes a central FL server capable of controlling the interaction items. One natural possible defense is to adopt a different FL architecture or protocol with no single centralized authority, targeting the core assumption of our attack. For example, in the peer-to-peer gossip-based decentralized recommendation model [30, 43], users exchange model weights/gradients and/or item interactions with other peers in the network, thus eliminating the need for a central server. Hierarchical FL is another example where users are grouped into clusters and communicate exclusively with some statically or dynamically determined “leaders” of the clusters [75, 42]. While these scenarios can prevent RAIFLE thanks to their network topology, they often assume that peers are honest-but-curious and thus can still be vulnerable to Sybil attacks where malicious peers collude to gain an oversized influence on the FL network and uncover user interaction data [12]. Additionally, even though the peers are also assumed to be able to communicate anonymously with one another, in reality, practical anonymous communication protocols often incur additional networking overhead (on top of the already increased communication expense from FL decentralization) and can still be subject to various deanonymization attacks [36]. That said, we believe these “serverless” FL schemes hold great potential to further protect user privacy, and we hope to see more development and adoption from both academia and industry.
A-D t-SNE Visualization of Gradients Under Manipulation
Appendix B Artifact Appendix
B-A Description & Requirements
B-A1 How to access
B-A2 Hardware dependencies
  • A commodity machine with at least 16GB of RAM and 30GB of storage.
  • Optional but highly recommended: A CUDA-capable NVIDIA GPU with at least 8GB of VRAM (preferably 12-16GB)
B-A3 Software dependencies
  • A (x86-64) Unix-based OS. (Windows WSL will probably also work, but might need some installation modifications.)
  • conda (such as Miniconda).
B-A4 Benchmarks
We evaluate on the following datasets:
  • Recommendation: MovieLens-100K, STEAM-200K
  • Learning-to-rank: MQ2008, MQ2007, MSLR-WEB10K
  • Image: ImageNet 2012 (validation split)
B-B Artifact Installation & Configuration
We use conda to set up the environment. Please see the README file in our repository for detailed instructions on how to set up as well as download the relevant datasets.
B-C Experiment Workflow
Please refer to the Evaluation section.
B-D Major Claims
  • (C1): RAIFLE achieves near-perfect AUC and significantly outperforms the state-of-the-art attack (IMIA) on federated recommendation (particularly the federated neural collaborative filtering algorithm), as demonstrated by experiment (E1) and reported in Tables V and VI of Section VI-A for MovieLens-100K and STEAM-200K.
  • (C2): RAIFLE outperforms traditional gradient inversion on federated online learning-to-rank, as demonstrated by experiments (E2) and (E3) and reported in Table VII for MQ2007 and MSLR-WEB10K (Section VI-B) and Table VIII for ImageNet (Section VI-C).
B-E Evaluation
All of our experiments are in IPython Jupyter Notebook inside the code folder. We recommend using Visual Studio Code to run our experiments. More detailed step-by-step instructions can be found in the README of our repository. For artifact evaluation, we make some suggested modifications to our code (e.g., omit some experiment configurations and reduce the number of simulations/users) to keep the evaluation under the time limit. Full-scale runs can be easily enabled by modifying our notebooks as instructed in the code comments.
Each experiment has the following structure:
  • The central FL server initializes a global FL model, performs manipulation on the FL training item features, then shares the model and items with users
  • Each user trains the global FL model locally using the manipulated data and their private interactions, applies differential privacy noise, then sends back the updated model weights to the server (directly or via secure aggregation).
  • The server inverts the gradient updates to guess the private interactions.
B-E1 Experiment (E1)
Federated Recommendation Systems: We run RAIFLE on the MovieLens-100K and STEAM-200K datasets with the Federated Neural Collaborative Filtering (FNCF) algorithm.
[Execution] Run all cells in experiment_rec.ipynb in order. The default dataset is MovieLens-100K. Cell 2 contains instructions on how to change the dataset. For artifact evaluation, we scaled down the number of users attacked to 30, which will take about  30 minutes to finish.
[…]
where {epsilon} refers to the local DP $\varepsilon$ parameter (inf means no DP). IMIA_0.0 means the IMIA defense is not applied, IMIA_1.0 means the IMIA defense is applied with L1 regularization factor 1.0.
Figure 10: Example output of the federated RS experiment on MovieLens-100K for only 3 users
B-E2 Experiment (E2)
Federated Online Learning to Rank (FOLTR) (non-image): We run RAIFLE on the MQ2007, MQ2008, and MSLR-WEB10K datasets with the Federated Pairwise Differentiable Gradient Descent (FPDGD) algorithm.
[Execution] Run cell 1, 2, and 3 in experiment_ltr.ipynb. Cell 2 contains instructions on how to change the dataset and other configs. For artifact evaluation, the default configuration is a linear ranker + a neural net ranker with 16 hidden units, MQ2007 dataset, and 16 queries per user (this will take about 5 hours, a full run on MQ2007 can take more than 1 day, MSLR-WEB10K is much longer).
[…]
where {model_name} is ‘linear_pdgd’, ‘neural_16_pdgd’, etc., {click_model_name} is either ‘informational’ or ‘navigational’, {num_query} is the number of queries per user (e.g., 16), {epsilon} is the local DP epsilon (inf means no DP), and {key} is either 0.0 or 1.0, where 0.0 means no manipulation and 1.0 means full manipulation.
Figure 11: Example output of FOLTR experiment on MQ2007 for only 3 users with the default configuration.
B-E3 Experiment (E3)
FOLTR with ImageNet: We run RAIFLE on the ImageNet dataset with a simple pointwise ranking algorithm. This is the only experiment where a GPU is recommended.
[Execution] Run cell 1, 2, 3, and 4 in experiment_ltr_cv.ipynb. The default configuration (scaled down for artifact evaluation) is ResNet18 as feature extractor, 30 rounds of simulation, and 5,000 images. Cell 2 contains instructions on how to change the feature extractor. Cell 3 generates the manipulated images. The batch size may need to be adjusted depending on how much GPU memory is available, e.g., 128 if 8GB, 256 if 12GB or more. Cell 3 can take 2-3 hours with this configuration. If a GPU is not available, we suggest using raifle_ltr_cv_colab.ipynb with Google Colab’s GPU runtime instead. Cell 4 simulates the attack and can take 2-3 hours (without GPU).
[…]
Figure 12: Example output of FOLTR experiment on ImageNet for only 3 simulation rounds with the default configuration.